# Boundary Constraint-free Biomechanical Model-Based Surface Matching for Intraoperative Liver Deformation Correction

### v1

@@ Abstract
In image-guided liver surgery, 3D-3D non-rigid registration methods play a crucial role in estimating the mapping between the preoperative model and the intraoperative surface represented as point clouds, addressing the challenge of tissue deformation. Typically, these methods incorporate a biomechanical model, represented as a finite element model (FEM), used to regularize a surface matching term. This paper introduces a novel 3D-3D non-rigid registration method. In contrast to the preceding techniques, our method uniquely incorporates the FEM within the surface matching term itself, ensuring that the estimated deformation maintains geometric consistency throughout the registration process. Additionally, we eliminate the need to determine zero-boundary conditions and applied force locations in the FEM. We achieve this by integrating soft springs into the stiffness matrix and allowing forces to be distributed across the entire liver surface. To further improve robustness, we introduce a regularization technique focused on the gradient of the force magnitudes. This regularization imposes spatial smoothness and helps prevent the overfitting of irregular noise in intraoperative data. Optimization is achieved through an accelerated proximal gradient algorithm, further enhanced by our proposed method for determining the optimal step size. Our method is evaluated and compared to both a learning-based method and a traditional method that features FEM regularization using data collected on our custom-developed phantom, as well as two publicly available datasets. Our method consistently outperforms or is comparable to the baseline techniques. Both the code and dataset will be made publicly available.
@@ I Introduction
To integrate subsurface information from preoperative CT/MRI into the surgeon’s intraoperative view, registration methods are employed to map critical information from the preoperative images into the intraoperative scene. Rigid registration is often used to align data collected in different coordinates. However, the initial configuration of the liver at the start of the intervention can significantly deviate from its preoperative state. Studies indicate that the deformation of the anterior surface of the liver can exceed 10 mm during laparoscopic liver surgery (LLS) and 7 mm during open liver surgery (OLS) [9]. In clinical contexts, the target registration error is expected to be below 5 mm [4]. Therefore, non-rigid registration becomes essential to correct for such deformations.
@@
This study focuses on the 3D-3D non-rigid registration problem in IGLS. Unlike the typical 3D-3D non-rigid registration problem in the computer vision [14], where the primary focus is achieving surface matching accuracy, the problem in IGLS emphasizes the necessity for the estimated organ deformation to be realistic, with a crucial consideration for the deformation beneath the tissue surface. Biomechanical models, specifically represented as finite element models (FEMs), are extensively employed in modeling tissue deformation [15] and are key to solving the 3D-3D non-rigid registration problem [4, 16, 5]. The biomechanical model not only constraints displacements to enforce realistic deformations but also provides insights into volumetric deformation, capturing the complexities of organ deformations beyond surface-level changes.
@@ I-A Related work
In deep learning-based techniques, such as neural networks, the biomechanical model is employed to simulate various deformations for training the neural network. Pfeiffer et al. [20] proposed V2Snet (volume-to-surface registration network), designed to estimate the deformation of a volume mesh to an intraoperative surface. Tagliabue et al. [21] introduced BA-Net (binary-attachment network) to predict the locations of the attachment points during tissue dissection. The predicted attachment points can be used as ZBCs to update a patient-specific biomechanical model intraoperatively [7].
Although utilizing the FEM as strain energy might restrict deformation, it fails to ensure coherent movement of volumetric vertices to maintain the original geometry. On the other hand, the direct solution of the FEM-based data term requires identifying Zero-Boundary Conditions (ZBCs) and force locations, as demonstrated in the existing method by Mestdagh et al. [16].
@@ I-B Contributions
In this paper, we introduce a novel 3D-3D non-rigid registration method that directly incorporates the FEM into the data term. Notably, our approach eliminates the need to manually identify ZBCs and force locations, setting it apart from existing methods. To account for the potential presence of noise associated with the reconstructed intraoperative point clouds, we integrate L2 regularization on the gradient of force magnitudes to ensure spatial smoothness. The optimization process involves an accelerated proximal gradient algorithm and a proposed optimal step decision to enhance acceleration and improve accuracy.
We validate our method using simulated and phantom datasets and two public datasets. Comparative assessments include validation against a learning-based learning-based approach and an optimization-based method utilizing FEM as regularization. The accuracy and robustness of our method relative to the benchmarks demonstrate its effectiveness.
@@ II Preliminary Information
During surgery, the observed intraoperative surface is represented by a point cloud $\mathbf{y}=\{\mathbf{y}_{1},...,\mathbf{y}_{m}|\mathbf{y}_{i}\in\mathbb{R}^{3}\}\in\mathbb{R}^{3m}$. The intraoperative surface can be obtained from various sources, such as CBCT, an RGBD depth camera, structure light, stereo laparoscope, or manual swabbing with an optically tracked probe.
The objective is to register a preoperative model with deformed intraoperative surface data to enable the mapping of the sub-surface anatomical structures identified in the preoperative images into the intraoperative scene. A preliminary alignment of the preoperative model with the intraoperative surface is necessary before the non-rigid step, typically achieved through a rigid transformation. Existing IGLS systems commonly perform this alignment manually [23, 24]; nevertheless, the manual alignment can further improve with a variant of the ICP [25, 26].
II-A Biomechinical model
@@
$\mathbf{K}\mathbf{u}=\mathbf{f},$ (1)
where $\mathbf{K}$ the $3n$ x $3n$ stiffness matrix determined by the FEM model of the liver geometry and material properties. In the forward problem, the displacement vector $\mathbf{u}$ throughout the domain $\mathbf{\Omega}$ can be determined if displacements and/or forces on the boundary $\mathbf{\partial\Omega}$ are known. The inverse registration problem solves for the boundary forces that generate a displacement field $\mathbf{u}^{*}$ that approximates the displacement field $\mathbf{u}$ between the partially observed intraoperative surface and preoperative model.
@@ II-B Stiffness Matrix and Boundary Conditions
The liver geometry and material properties are encoded in the stiffness matrix of the finite element model. This matrix characterizes the object’s resistance to nodal deformation when subjected to nodal forces. In the case of a linearly elastic biomechanical model, [27], Young’s modulus ($E$), and Poisson’s ratio ($\nu$) account for the material’s elastic properties.
@@ III Methods
Figure 1: Overview of the proposed method. The process takes a biomechanical model generated from the preoperative volume model and an intraoperative point cloud. Force optimization leads to volumetric deformation, aligning the preoperative volume model with the intraoperative point cloud.
Our method employs a forces-driven approach to deform the source model for optimal alignment with the target point cloud. Fig. 1 provides an overview of our registration method. The objective function comprises a surface matching term and an L2 regularization term. The surface matching term utilizes a correspondence matrix generated by a closest point operator and a FEM. Minimization of the objective function is achieved through an accelerated proximal gradient algorithm. During the force optimization phase, we initially determine the optimal step for calculating Nesterov’s accelerated gradient [28]. Subsequently, the gradient is employed to update forces using a proximal point operator. In each iteration, we first update the correspondence matrix via the closest point operator, then update the force through force optimization; finally, we update the displacements using the FEM. After a set number of iterations, the predicted volumetric deformation enables the preoperative volume model to align with the intraoperative point cloud. In the following section, we will detail each of these steps.
@@ III-A Objective function
We perform the registration by determining a surface force vector $\mathbf{f}$ which results in a displacement vector $\mathbf{u_{f}}$ that minimizes the objective function:
$\min_{\mathbf{f}\in\mathbf{F}}\Phi(\mathbf{f})=\beta R(\mathbf{f})+J(\mathbf{u_{f}}),$ (2)
where $J$ is a data term that measures the discrepancy between the predicted deformed preoperative surface $\mathbf{\partial\Omega}_{\mathbf{u}}$ and the observed intraoperative point cloud data, $R$ is a regularization term that enforces a given property on the forces to prevent overfitting of the data term, $\beta$ is a weight between the data and regularization term, and $\mathbf{F}$ is the set of force vectors containing prior information about the location of nonzero forces on the liver’s surface.
Data Term. The data term is given by
$J(\mathbf{u})=\frac{1}{2}\|\mathbf{C}(\mathbf{x}+\mathbf{u})-\mathbf{y}\|^{2},$ (3)
where $\mathbf{C}\in\mathbb{R}^{3m\times 3n}$ is a soft correspondence matrix, $\mathbf{x}\in\mathbb{R}^{3n}$ are spatial location of the preoperative mesh nodes, $\mathbf{u}\in\mathbb{R}^{3n}$ are displacements associated with the nodes, and $\mathbf{y}\in\mathbb{R}^{3m}$ is the intraoperative point cloud. The $\mathbf{C}$ will be detailed in Sec. III-B.
The liver biomechanical model can be incorporated into Eq. 3 by using the relationship between $\mathbf{u}$ and $\mathbf{f}$ given by Eq. 1:
$J(\mathbf{f})=\frac{1}{2}\|\mathbf{C}(\mathbf{x}+\mathbf{K}^{-1}\mathbf{f})-\mathbf{y}\|^{2}.$ (4)
The resulting data term now becomes a function of the forces applied to the surface of the liver.
Regularization Term. We introduce Tikhonov (L2) regularization to enforce the spatial smoothness of the surface forces:
$R(\mathbf{f})=\frac{1}{2}\|\nabla_{w}\mathbf{f}\|^{2}=\frac{1}{2}\sum_{i=1}^{n_{s}}\sum_{j\in{N}(i)}(\mathbf{f}_{j}-\mathbf{f}_{i})^{2}w_{ij}.$ (5)
where $\mathbf{f}_{i}$ is the force vector at node $i$, ${N}(i)$ represents the $j$ nodes connected to node $i$, and the weight function $w_{ij}=\frac{1}{d_{ij}}$. $d_{ij}$ represents the Euclidean distance between node $i$ and $j$.
@@ III-B Correspondence Matrix
The correspondence matrix $\mathbf{C}$ contains the correspondences between the points in the intraoperative point cloud and preoperative model. Ideally, if corresponding fiducial points or anatomical landmarks could be consistently identified in pre- and intraoperative spaces, the correspondence matrix would be a one-to-one binary matrix. Obtaining an accurate binary correspondence matrix is nontrivial unless manual annotation is used.
Instead of an accurate binary correspondence matrix, a soft correspondence matrix is often constructed, indicating the probabilistic correspondence between a point in one space and another. The elements $C_{ij}$ of the correspondence matrix $\mathbf{C}$ denote the probability that the intraoperative point $\mathbf{y}_{i}$ corresponds to the preoperative mesh point $\mathbf{x}_{j}+\mathbf{u}_{j}$. Each row of the correspondence matrix should sum to one.
@@
Here we use the closet point operator to construct a soft correspondence matrix. For a given intraoperative point $\mathbf{y}_{i}$, the closest point operator determines its closest point $\tilde{\mathbf{y}}_{i}$:
@@
$\lambda_{i,j,k}$ sums to one, representing confidence values $C_{i,(i,j,k)}$. Thus, the 3m x 3n correspondence matrix is extremely sparse, where each row contains only three non-zero entries corresponding to the barycentric coordinates. The closest point operator and the resulting correspondence matrix are a function of the undeformed mesh $\mathbf{x}$, intraoperative point cloud $\mathbf{y}$, and the nodal displacements $\mathbf{u}$. The correspondence matrix is recalculated during the iterative registration using the current estimate of the displacements.
@@ III-C Optimization
Nesterov’s accelerated gradient [28] and a proximal point operator [31] are used to minimize the objective function Eq. 2. The following steps
@@
$\overline{\mathbf{f}}^{k}=\mathbf{p}^{k}-\alpha^{k}\nabla J(\mathbf{p}^{k}),$ (8)
$\mathbf{f}^{k+1}=\arg\min_{\mathbf{f}}\Bigl{(}\frac{\beta}{2}\|\nabla_{w}\mathbf{f}\|^{2}+\frac{1}{2\alpha^{k}}\|\mathbf{f}-\overline{\mathbf{f}}^{k}\|^{2}\Bigr{)},$ (9)
@@
Nesterov’s accelerated gradient is given by Eq. 7 and Eq. 8. $\mathbf{p}^{k}$ is a momentum term that carries momentum from the previous iteration. Eq. 8 updates the forces, where $\nabla J(\mathbf{p}^{k})$ is the gradient of the data term (Eq. 4) with respect to $\mathbf{p}$, and $\alpha^{k}$ is the step size. The $\nabla J(\mathbf{p}^{k})$ is given by
$\begin{split}\nabla J(\mathbf{p}^{k})=\Bigl{(}(\mathbf{x}+\mathbf{K}^{-1}\mathbf{p}^{k})^{T}\frac{\partial\mathbf{C}^{T}}{\partial\mathbf{f}}+\mathbf{K}^{-1}\mathbf{C}^{T}\Bigr{)}\cdot\\ \Bigl{(}\mathbf{C}(\mathbf{x}+\mathbf{K}^{-1}\mathbf{p}^{k})-\mathbf{y}\Bigr{)},\end{split}$ (10)
where $\frac{\partial\mathbf{C}^{T}}{\partial\mathbf{f}}$ is a 3-dimensional array (3 tenors). It yields a $3n$ x $3m$ matrix when multiplied by the vector $(\mathbf{x}+\mathbf{K}^{-1}\mathbf{p})^{T}$
@@
$\alpha^{k}=\arg\min_{\alpha}\frac{1}{2}\Bigl{\|}\mathbf{C}\Bigl{(}\mathbf{x}+\mathbf{K}^{-1}(\mathbf{p}^{k}-\alpha\nabla J(\mathbf{p}^{k}))\Bigr{)}-\mathbf{y}\Bigr{\|}^{2},$ (11)
@@
$\alpha^{k}=\frac{\Bigl{(}\mathbf{C}\mathbf{K}^{-1}\nabla J(\mathbf{p}^{k})\Bigr{)}^{T}\Bigl{(}\mathbf{C}(\mathbf{x}+\mathbf{K}^{-1}\mathbf{p}^{k})-\mathbf{y}\Bigr{)}}{\Bigl{(}\mathbf{C}\mathbf{K}^{-1}\nabla J(\mathbf{p}^{k})\Bigr{)}^{T}\Bigl{(}\mathbf{C}\mathbf{K}^{-1}\nabla J(\mathbf{p}^{k})\Bigr{)}}.$ (12)
Proximal point operator, given by the Eq. 9, enforces the force vector $\mathbf{f}^{k+1}$ to be smooth, which is the result of the trade-off between minimizing the function $\|\nabla_{w}\mathbf{f}\|^{2}$ and its proximity to $\overline{\mathbf{f}}^{k+1}$.
To solve the proximal equation Eq. 9, an approximate solution is obtained by employing an iterative procedure. The minimization of proximal equation Eq. 9 yields the following equation:
$(\mathbf{I}+\beta\alpha^{k}\Delta_{w})\mathbf{f}^{k+1}=\overline{\mathbf{f}}^{k},$ (13)
where the graph Laplacian $\Delta_{w}\mathbf{f}_{i}$ is given by
$\Delta_{w}\mathbf{f}_{i}=\sum_{j\in{N}(i)}(\mathbf{f}_{i}-\mathbf{f}_{j})w_{ij},$ (14)
The approximate solution of the previous equation is obtained by employing a Gauss–Seidel iterative procedure [32] that iteratively updates the unknown component of $\mathbf{f}_{i}^{k+1}$, then subsequently used to update the remaining components. The solution can be written component-wise as:
$\mathbf{f}_{i}^{k+1}=\frac{\overline{\mathbf{f}}_{i}^{k}+\beta\alpha^{k}\sum_{j\in N(i)}\mathbf{f}_{j}w_{ij}}{1+\beta\alpha^{k}\sum_{j\in N(i)}w_{ij}}.$ (15)
The $\mathbf{C}$ should also be updated with forces being updated. We employ a lagged method, utilizing the nodal displacements $\mathbf{u}^{k}$ from the previous iteration, to update the correspondence matrix in each iteration.

### v2

@@ Abstract
In image-guided liver surgery, 3D-3D non-rigid registration methods play a crucial role in estimating the mapping between the preoperative model and the intraoperative surface represented as point clouds, addressing the challenge of tissue deformation. Typically, these methods incorporate a biomechanical model, represented as a finite element model (FEM), used to regularize a surface matching term. This paper introduces a unique 3D-3D non-rigid registration method, which, in contrast to the preceding techniques. Our method incorporates the FEM within the surface-matching term itself, inherently ensuring that the estimated deformation maintains geometric consistency throughout the registration process. Additionally, we eliminate the need to identify zero-boundary conditions and applied force locations in the FEM. We achieve this by integrating soft springs into the stiffness matrix and allowing forces to be distributed across the entire liver surface. Optimization is achieved through an accelerated gradient algorithm, further enhanced by our proposed method for determining the optimal step size. We evaluated our method and compared it to several state-of-the-art methods across various datasets. Our straightforward and effective approach consistently outperformed or achieved comparable performance to the state-of-the-art methods. We will make both the code and dataset publicly available.
@@ I Introduction
To integrate subsurface information from preoperative CT/MRI into the surgeon’s intraoperative view, registration methods are employed to map critical information from the preoperative images into the intraoperative scene. Rigid registration is often used to align data collected in different coordinates. However, the initial configuration of the liver at the start of the intervention can significantly deviate from its preoperative state. Studies indicate that the deformation of the anterior surface of the liver can exceed 10 mm during laparoscopic liver surgery (LLS) and 7 mm during open liver surgery (OLS) [9]. In clinical contexts, the target registration error is expected to be below 5 mm [4]. Therefore, non-rigid registration may become essential to correct for such deformations.
@@
This study focuses on the 3D-3D non-rigid registration problem in IGLS. Unlike the typical 3D-3D non-rigid registration problem in the computer vision [14], where the primary focus is achieving surface matching accuracy, the problem in IGLS emphasizes the necessity for the estimated organ deformation to be realistic, with a crucial consideration for the deformation beneath the tissue surface. Biomechanical models, specifically represented as finite element models (FEMs), have been extensively employed in modeling tissue deformation [15] and are key to solving the 3D-3D non-rigid registration problem [4, 16, 5]. The biomechanical model not only constraints displacements to enforce realistic deformations but also provides insights into volumetric deformation, capturing the complexities of organ deformations beyond surface-level changes.
@@ I-A Related work
Despite being in the era of deep learning, learning-based methods for liver deformation registration are still constrained by several factors: the scarcity of large datasets for training and testing, issues with interpretability, and suboptimal accuracy. Additionally, effectively combining learning-based methods with biomechanical model-based approaches to address registration problems still remains a challenge. In learning-based techniques, the biomechanical model is often employed to simulate various deformations for training the neural network. Pfeiffer et al. [20] proposed V2Snet (volume-to-surface registration network), designed to estimate the deformation of a volume mesh to an intraoperative surface. Tagliabue et al. [21] introduced BA-Net (binary-attachment network) to predict the locations of the attachment points during tissue dissection. The predicted attachment points can be used as ZBCs to update a patient-specific biomechanical model intraoperatively [7].
Although utilizing the FEM as strain energy might restrict deformation, it fails to ensure coherent movement of volumetric vertices to maintain the original geometry. On the other hand, the direct solution of the FEM-based data term requires the identification of Zero-Boundary Conditions (ZBCs) and force locations, as demonstrated in the existing method by Mestdagh et al. [16].
@@ I-B Contributions
Most non-rigid registration methods incorporate the FEM as the strain energy/regularization term. Only few methods include the FEM as the surface matching data term, which requires the boundary conditions to be manually identified. To mitigate the inconvenience of the manual intervention, which is also prone to error, our proposed method incorporates the modified FEM into the data term using a modified stiffness matrix. This approach stabilizes the stiffness matrix and eliminates the need for manual identification of ZBCs and fore locations, hence rendering our method unique among the existing state-of-the-art methods.
To solve our formulation, we employ the Nesterov accelerated gradient algorithm and derive the necessary gradients. Additionally, we propose a novel optimal step size strategy that eliminates the need for manual step size tuning.
We validate our method using our both simulated and experimentally collected phantom datasets, as well as two publicly available datasets. Comparative assessments are conducted against two open-source methods - a learning-based approach and an optimization-based method utilizing FEM for regularization, along with state-of-the-art closed-source methods. This straightforward yet effective approach consistently demonstrates its effectiveness relative to the benchmarks.
@@ II Preliminary Information
[…]
The objective is to register a preoperative model with the deformed intraoperative surface data to map the sub-surface anatomical structures identified in the preoperative images into the intraoperative scene. A preliminary alignment of the preoperative model with the intraoperative surface is necessary before the non-rigid step, typically achieved through a rigid transformation. Existing IGLS systems commonly perform this alignment manually [24, 25]; nevertheless, the manual alignment can be further improved with a variant of the ICP [26, 27].
II-A Biomechanical Model
@@
$\mathbf{P}(\mathbf{u})=\mathbf{f},$ (1)
where $\mathbf{P(u)}=\{P_{1}(\mathbf{u}),...,P_{n}(\mathbf{u})\}$ is a vector of nonlinear functions of $\mathbf{u}$ and are determined by the FEM model of the liver geometry and material properties. For a liver modeled by a linear elastic material, the forces are now a linear function of $\mathbf{u}$ where $\mathbf{P}(\mathbf{u})=\mathbf{K}\mathbf{u}$ with $\mathbf{K}$ is a constant $3n$ x $3n$ stiffness matrix. For a nonlinear hyperelastic material, $\mathbf{P}(\mathbf{u})$ is a nonlinear function of the displacements u. The tangent stiffness matrix $\mathbf{K}_{T}$ defined as the derivative of $\mathbf{P}(\mathbf{u})$ with respect $\mathbf{u}$
$\frac{\partial\mathbf{f}}{\partial\mathbf{u}}=\frac{\partial\mathbf{P}(\mathbf{u})}{\partial\mathbf{u}}=\mathbf{K}_{T}(\mathbf{u}),$ (2)
is a matrix used in each iteration of the solution of a nonlinear problem, and it changes with each iteration. For a linear elastic material, the tangent stiffness matrix is simply the stiffness matrix.
In the forward problem, the displacement vector $\mathbf{u}$ throughout the domain $\mathbf{\Omega}$ can be determined if displacements and/or forces on the boundary $\mathbf{\partial\Omega}$ are known. The inverse registration problem solves for the boundary forces that generate a displacement field $\mathbf{u}^{*}$ that approximates the displacement field $\mathbf{u}$ between the partially observed intraoperative surface and preoperative model.
@@ II-B Stiffness Matrix and Boundary Conditions
The liver geometry and material properties are encoded in the stiffness matrix of the finite element model. This matrix characterizes the object’s resistance to nodal deformation when subjected to nodal forces. In the case of a linearly elastic biomechanical model[28], Young’s modulus ($E$), and Poisson’s ratio ($\nu$) account for the material’s elastic properties.
@@ III Methods
Figure 1: Overview of the proposed method. The process takes a biomechanical model generated from the volumetric liver model (typically segmented from a CT or MRI image) and an intraoperative point cloud. Force optimization leads to volumetric deformation, aligning the preoperative volume model with the intraoperative point cloud.
Our method employs a force-driven approach to deform the source model for optimal alignment with the target point cloud. Fig. 1 provides an overview of our registration method. The objective function comprises a surface matching term. The surface matching term utilizes a correspondence matrix generated by a closest point operator and a FEM. Minimization of the objective function is achieved through an accelerated gradient algorithm. During the force optimization phase, we initially determine the optimal step for calculating Nesterov’s accelerated gradient [29]. Subsequently, the gradient is employed to update forces. In each iteration, we first update the correspondence matrix via the closest point operator, then update the force through force optimization; finally, we update the displacements using the FEM. After a set number of iterations, the predicted volumetric deformation enables the preoperative volume model to align with the intraoperative point cloud. In the following section, we will detail each of these steps.
@@ III-A Objective Function
We perform the registration by determining a surface force vector $\mathbf{f}$ that results in a displacement vector $\mathbf{u_{f}}$, which minimizes the objective function:
$\min_{\mathbf{f}\in\mathbf{F}}\Phi(\mathbf{f})=J(\mathbf{u_{f}}),$ (3)
where $J$ is a data term that measures the discrepancy between the predicted deformed preoperative surface $\mathbf{\partial\Omega}_{\mathbf{u}}$ and the observed intraoperative point cloud data.
The data term is given by
$J(\mathbf{u})=\frac{1}{2}\|\mathbf{C{\color[rgb]{0,0,0}(u)}}(\mathbf{x}+\mathbf{u})-\mathbf{y}\|^{2},$ (4)
where $\mathbf{C}\in\mathbb{R}^{3m\times 3n}$ is a soft correspondence matrix, $\mathbf{x}\in\mathbb{R}^{3n}$ are spatial location of the preoperative mesh nodes, $\mathbf{u}\in\mathbb{R}^{3n}$ are displacements associated with the nodes, and $\mathbf{y}\in\mathbb{R}^{3m}$ is the intraoperative point cloud. The correspondence matrix $\mathbf{C}$ maps the deformed to a point on the surface will be detailed in Sec. III-B. Our data term is a reformulation of the data term from Mestdagh et al. [16, 19] where we have explicitly incorporated the correspondence matrix and the dependence of $J(\mathbf{u})$ on $\mathbf{u}$.
@@ III-B Correspondence Matrix
The correspondence matrix $\mathbf{C}$ contains the correspondences between the points in the intraoperative point cloud and preoperative model. Ideally, if corresponding fiducial points or anatomical landmarks could be consistently identified in the pre- and intraoperative spaces, the correspondence matrix would be a one-to-one binary matrix. However, obtaining an accurate binary correspondence matrix is non-trivial unless manual annotation is used, which is highly user-dependent and prone to error.
Instead of a binary correspondence matrix, a soft correspondence matrix is often constructed, indicating the probabilistic correspondence between a point in one space and another. The elements $C_{ij}$ of the correspondence matrix $\mathbf{C}$ denote the probability that the intraoperative point $\mathbf{y}_{i}$ corresponds to the preoperative mesh point $\mathbf{x}_{j}+\mathbf{u}_{j}$. Each row of the correspondence matrix should sum to one.
@@
Here, we use the closest point operator to construct a soft correspondence matrix. For a given intraoperative point $\mathbf{y}_{i}$, the closest point operator determines its closest point $\tilde{\mathbf{y}}_{i}$:
@@
The $\lambda_{i,j,k}$ sums to one, representing confidence values $C_{i,(i,j,k)}$. Thus, the 3m x 3n correspondence matrix is extremely sparse, where each row contains only three non-zero entries corresponding to the barycentric coordinates. The closest point operator and the resulting correspondence matrix are a function of the undeformed mesh $\mathbf{x}$, intraoperative point cloud $\mathbf{y}$, and the nodal displacements $\mathbf{u}$.
@@ III-C Optimization
Nesterov’s accelerated gradient [29] is used to minimize the objective function Eq. 3. Nesterov’s accelerated gradient is a momentum-based optimizer that adds a fraction of the previous update to the current one, creating a momentum effect that accelerates the algorithm toward the minimum. The following steps
@@
$\mathbf{f}^{k+1}=\mathbf{p}^{k}-\alpha^{k}\nabla J(\mathbf{p}^{k}),$ (7)
@@
Nesterov’s accelerated gradient is given by Eq. 6 and Eq. 7. $\mathbf{p}^{k}$ carries momentum from the previous iteration by combining the force estimates of the current and previous iterations. Eq. 7 updates the forces, where $\nabla J(\mathbf{p}^{k})$ is the gradient of the data term (Eq. 4) with respect to $\mathbf{p}$, and $\alpha^{k}$ is the step size. Using the relationship between $\mathbf{u}$ and $\mathbf{f}$ given by Eq.2, the $\nabla J(\mathbf{p}^{k})$ is given by
${\color[rgb]{0,0,0}\nabla J(\mathbf{p}^{k})=\mathbf{K}_{T}^{-1}\Bigl{(}(\mathbf{x}+\mathbf{u}^{k})^{T}\frac{\partial\mathbf{C}^{T}}{\partial\mathbf{u}}+\\ \mathbf{C}^{T}(\mathbf{C}(\mathbf{x}+\mathbf{u}^{k}-\mathbf{y})\Bigr{)}},$ (8)
where for a linear elastic material $\mathbf{K}_{T}$ is a constant and for a nonlinear material $\mathbf{K}_{T}$ is a function of $\mathbf{u}^{k}$ and can be updated as needed. The correspondence matrix is recalculated during the iterative registration using the current estimate of the displacements. To simplify the calculation of $\nabla J(\mathbf{p}^{k})$ we employ an iteration-lagging method which updates the correspondence matrix for each iteration using nodal displacements $\mathbf{u}^{k}$ computed from the previous iteration. The correspondence matrix is now considered fixed and its variation with respect to $\mathbf{u}$ can be neglected (i.e. $\frac{\partial\mathbf{C}^{T}}{\partial\mathbf{u}}=0$). Now $\nabla J(\mathbf{p}^{k})$ is given by
$\nabla J(\mathbf{p}^{k})=\mathbf{K}_{T}^{-1}\mathbf{C}^{T}\Bigl{(}\mathbf{C}(\mathbf{x}+\mathbf{u}^{k})-\mathbf{y}\Bigr{)}.$ (9)
@@
$\alpha^{k}=\arg\min_{\alpha}\frac{1}{2}\Bigl{\|}\mathbf{C}\Bigl{(}\mathbf{x}+\mathbf{K}^{-1}_{T}(\mathbf{p}^{k}-\alpha\nabla J(\mathbf{p}^{k}))\Bigr{)}-\mathbf{y}\Bigr{\|}^{2},$ (10)
@@
$\alpha^{k}=\frac{\Bigl{(}\mathbf{C}\mathbf{K}^{-1}_{T}\nabla J(\mathbf{p}^{k})\Bigr{)}^{T}\Bigl{(}\mathbf{C}(\mathbf{x}+\mathbf{K}^{-1}_{T}\mathbf{p}^{k})-\mathbf{y}\Bigr{)}}{\Bigl{(}\mathbf{C}\mathbf{K}^{-1}_{T}\nabla J(\mathbf{p}^{k})\Bigr{)}^{T}\Bigl{(}\mathbf{C}\mathbf{K}^{-1}_{T}\nabla J(\mathbf{p}^{k})\Bigr{)}}.$ (11)